OpenStudy (anonymous):

**Calc 2 Help Needed** Solve the initial value problem below: (t^2 + 2t)dx/dt = 2x + 5 with t, x>0 and x(1) = 1

OpenStudy (anonymous):

Rewrite as \[\frac{dx}{2x+5}=\frac{dt}{t^2+2t}\] \[\frac{dx}{2x+5}=\frac{dt}{t^2+2t+1-1}\] \[\frac{dx}{2x+5}=\frac{dt}{(t+1)^2-1}\] Integrate both sides

OpenStudy (anonymous):

ok lemme see.

OpenStudy (anonymous):

ok so now.. ln(2x + 5) = ln(t) - ln(t + 2)

OpenStudy (anonymous):

that sound right?